OpenStudy (anonymous):

An equation is shown below: km2 - c = n Which equation is solved for m?

OpenStudy (nirmalnema):

km2 - c = n km^2=n+c(adding c both side) m^2=(n+c)/k (dividing by k both side) now taking root both side m=rt. ov (n+c)/k